OpenStudy (anonymous):

Can someone help me out with this word problem? The residents of a city voted on whether to raise property taxes. The ratio of yes votes to no votes was 7 to 4. If there were 7843 total votes, how many yes votes were there?

OpenStudy (anonymous):

7y +4n = 7843

OpenStudy (danjs):

y/n = 7/4 AND y+n = 7843

OpenStudy (danjs):

2 equations and 2 unknowns

OpenStudy (danjs):

the ratio of yes to no is 7 to 4 y/n = 7/4 the total votes n+y is 7843 n + y = 7843

OpenStudy (danjs):

solve